OpenStudy (anonymous):

What is the vertex and axis of symmetry for y=2x^+6x+5

Parth (parthkohli):

For a parabola, do you know what the vertex is?

Parth (parthkohli):

I mean there's a formula to find the coordinates of the vertex. Are you aware of that?

OpenStudy (anonymous):

The vertex is (-3/2,1/2)

OpenStudy (luigi0210):

Find the vertex using -\(\large \frac{b}{2a}\) for the x-coordinate. Plug it back in to get y.

Parth (parthkohli):

Yes, that is very much it. For the second part, do you see how the axis of symmetry of a parabola is?

Parth (parthkohli):

It's basically a vertical line passing through the vertex. So it'd have \(x = -\dfrac{b}{2a}\) everywhere.

Parth (parthkohli):

@linda3 please help!

OpenStudy (linda3):

I think you got it right @ParthKohli

OpenStudy (luigi0210):

Check it using calc as well.. for us. \(\frac{d}{dx}2x^2+6x+5=4x+6\) So the vertex is at \(x=\frac{-3}{2}\) like they said.

OpenStudy (anonymous):

the function is not symmetric with respect to the y-axis or the origin.